5. Hydrogen undergoes a polytropic process with n=0.7 from an initial specific volume of 22 ft3/lb and initial temperature of 200° F to a final specific volume of 16 ft3/lb. The entropy decrease during the process is 7.5 Btu/°R. Model the hydrogen as an ideal gas with constant specific heats. Evaluate the specific heat at 200° F. Determine the mass of the hydrogen.
Answer #1

V₂ = 16 ft3/lb 1= 22 +t%e6 Ti= 200°F = 200 +459. 67 Ti= 659.67 R AS = - 7.5 letulor n=0.7 solun For procues 1-2 ♡ no 0.7-1 T2